A source of light O is located at a distance of 8cm from a plane mirror. The reflected ray O’  is detected 2 cm from the mirror at a vertical displacement of 12 cm. The point of reflection is at a distance of:
a)6 cm from X
b)7.2 cm from X
c)2.4 cm from Y
d)3.6 cm from Y

Grade:8

1 Answers

Ravleen Kaur
6 years ago
Hello Student,

In the diagram angle i = angle r (by the law of reflection)
Sin i = 12-x/8
Sin r =x/2
Sin i = Sin r
12-x/8 = x/2
x = 12/5 = 2.4 cm from Y

Regards,
IHE